# The reflection complexity of sequences over finite alphabets

### v1

@@ 2 General results
Given a sequence $\mathbf{x}$, we can decompose its factor complexity function $\rho_{\mathbf{x}}$ and its reflection complexity function $r_{\mathbf{x}}$ by using the following functions:
@@
    The number $\operatorname{Unr}_{\mathbf{x}}(n)$ of “unreflected” length-$n$ factors $w$ of $\mathbf{x}$ such that $w^{R}$ is not a factor of $\mathbf{x}$;
@@
    The number $\operatorname{Ref}_{\mathbf{x}}(n)$ of “reflected” length-$n$ factors $w$ of $\mathbf{x}$ such that $w^{R}$ is also a factor of $\mathbf{x}$; and
@@
    The number $\operatorname{Pal}_{\mathbf{x}}(n)$ of length-$n$ palindrome factors $w$ of $\mathbf{x}$ (i.e., the palindrome complexity function of $\mathbf{x}$, see [4]).
@@
$\begin{array}[]{l}\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)+\operatorname{Ref}_{\mathbf{x}}(n),\\ r_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)+\dfrac{1}{2}(\operatorname{Ref}_{\mathbf{x}}(n)-\operatorname{Pal}_{\mathbf{x}}(n))+\operatorname{Pal}_{\mathbf{x}}(n).\end{array}$ (5)
@@
Let $\mathbf{f}=01101010\cdots$ denote the Fibonacci sequence, which is the fixed point of $0\mapsto 01,1\mapsto 0$. Its length-$5$ factors are $u_{1}=01001$, $u_{2}=10010$, $u_{3}=00101$, $u_{4}=01010$, $u_{5}=10100$, and $u_{6}=00100$. Observe that $u_{4}$ is an unreflected factor (Type 1 above), $u_{1},u_{2},u_{3},u_{5}$ are reflected (Type 2), and $u_{6}$ is a palindrome (Types 2 and 3). We obtain $r_{\mathbf{f}}(6)=1+\frac{1}{2}(5-1)+1=4$.
There is a rich interplay among the complexity functions $\operatorname{Unr}_{\mathbf{x}}$, $\operatorname{Ref}_{\mathbf{x}}$, $\rho_{\mathbf{x}}$, $r_{\mathbf{x}}$, and $\operatorname{Pal}_{\mathbf{x}}$, which motivates the study of the combinations of these functions indicated in Equalities (5). This is illustrated below.
@@
This lemma implies the following bounds on the ratio $\frac{r}{\rho}$.
@@
For any sequence $\mathbf{x}$ and for all $n\geq 0$, we have
@@
Furthermore, the equality cases are as follows.
@@
It is known that if a sequence $\mathbf{x}$ is reversal-closed, then $\mathbf{x}$ is recurrent: it suffices to adapt the proof of [40, Proposition 1, p. 176], as indicated in [22]. Also note that if a sequence $\mathbf{x}$ is uniformly recurrent and contains infinitely many distinct palindromes, then $\mathbf{x}$ is reversal-closed [15, Theorem 3.2].
@@
If $\mathbf{x}$ is reversal-closed, then $r_{\mathbf{x}}=\frac{1}{2}(\rho_{\mathbf{x}}+\operatorname{Pal}_{\mathbf{x}})$ from Theorem 9(b) above. Now suppose that $\mathbf{x}$ has an unreflected factor $w$. Since $\mathbf{x}$ is uniformly recurrent, any long enough factor of $\mathbf{x}$ contains $w$ as a factor, which implies that this long factor itself is unreflected. This exactly says that $\operatorname{Ref}_{\mathbf{x}}(n)=0$ for $n$ large enough (and in particular $\operatorname{Pal}_{\mathbf{x}}(n)=0$ for $n$ large enough). This implies from Equalities (5) that, for $n$ large enough, $\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)$, and so $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$. ∎
@@
Consider the sequence $\mathbf{x}$ on $\{0,1\}$ whose $n$th prefix $x_{n}$ is given recursively as follows: $x_{0}=01$ and $x_{n+1}=x_{n}01x_{n}^{R}$ for all $n\geq 0$. See [19, Section 3] or [15, Example 3.1]. The sequence $\mathbf{x}$ is uniformly recurrent, reversal-closed, and $2$-automatic and accepted by a DFAO of 6 states (see, e.g., [6]), and contains only a finite number of palindromes. Furthermore, for all sufficiently large $n$, we have $r_{\mathbf{x}}(n)=\dfrac{1}{2}\rho_{\mathbf{x}}(n)$.
@@
It is also possible to construct an aperiodic automatic sequence where the only palindromes are of length 1, but there are reflected factors of each length $>1$. In this regard, we let $g_{n}$ be the prefix of length $22^{n}-$ 2 of $(012)^{\omega}$. Then an example of an automatic sequence satisfying the desired properties is
@@
where [74, Theorem 1] is required (observe that we intertwine the sequences $(3456)^{\omega}$ and $g_{1}g_{2}^{R}g_{3}g_{4}^{R}\cdots$ to build our sequence).
@@
We also provide a construction of an automatic sequence $\mathbf{x}$ such that $r_{\mathbf{x}}(n+1)<r_{\mathbf{x}}(n)$ for all odd $n\geq 3$. In particular, let $\mathbf{x}$ denote the sequence given by applying the coding $a,b,d\mapsto 1$ and $c\mapsto 0$ to the fixed point, starting with $a$, of the morphism defined by $a\mapsto ab$, $b\mapsto cd$, $c\mapsto cd$, and $d\mapsto bb$. This gives us sequence [85, A039982] in the OEIS. Computing the reflection complexity of $\mathbf{x}$ (e.g., using Walnut) gives that
@@
Actually we even have, for this sequence $\mathbf{x}$, that $r_{\mathbf{x}}(n+1)=r_{\mathbf{x}}(n)-1$ for all odd $n\geq 3$.
With extra hypotheses on a sequence $\mathbf{x}$, we can give more precise results in comparing the respective growths of reflection and factor complexities. We will need Theorem 17 below. Note that Part (b) of this theorem was originally stated for uniformly recurrent sequences: see [14, Theorem 1.2]. However, its proof only requires the sequences to be recurrent (see [15, p. 449] and also [23, Footnote, p. 493]). Furthermore we have seen that a reversal-closed sequence must be recurrent (see Remark 10). Thus we can state the theorem as follows (also see Theorem 11).
@@
We deduce the following theorem from Theorem 17.
@@
Using Theorem 9(a) and (b), Remark 10 and Theorem 17(b), we have
@@
for all $n\geq 0$. Thus
$\frac{1}{2}\rho_{\mathbf{x}}(n)\leq r_{\mathbf{x}}(n)\leq\frac{1}{2}\rho_{\mathbf{x}}(n+1)+1$
for all $n\geq 0$, as desired. ∎
@@
for all $n\geq 1$. The statement follows from this and Theorem 9(a) and (b). ∎
@@
when $n$ tends to infinity. In particular, this equivalence holds if $\mathbf{x}$ is non-eventually periodic, reversal-closed and morphic.
@@
Let $\mathbf{x}$ be a non-eventually periodic and reversal-closed sequence. If $\rho_{\mathbf{x}}(n+1)\sim\rho_{\mathbf{x}}(n)$, then, from Theorem 19, we obtain that $r_{\mathbf{x}}(n)\sim\frac{1}{2}\rho_{\mathbf{x}}(n)$ when $n$ tends to infinity. Now, if $\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}=o(n)$, we obtain, from Theorems 9 and 20, and using the fact that $\rho_{\mathbf{x}}$ is non-decreasing,
@@
Now suppose that, in addition, the sequence $\mathbf{x}$ is morphic. We know that either $\rho_{\mathbf{x}}(n)=\Theta(n^{2})$ or $\rho_{\mathbf{x}}(n)=O(n^{3/2})$ (see [41]). In the first case, then $\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}$ is bounded, and hence $o(n)$. If $\rho_{\mathbf{x}}(n)=O(n^{3/2})$, since ${\mathbf{x}}$ is not eventually periodic (hence $\rho_{\mathbf{x}}(n)\geq n+1$), we have
@@
The upper bound in Theorem 20 raises questions as to growth properties of the function $r_{\mathbf{x}}$ more generally, apart from the case where the set of factors of $\mathbf{x}$ satisfies the hypotheses of Theorem 20. This leads us toward the growth property in Theorem 22 below.
@@
For a factor $w$ of $\mathbf{y}$, define $[w]_{1}=1$ if $w$ is a palindrome, and 00 otherwise. Similarly, $[w]_{2}=1$ if $w^{R}$ is not a factor of $\mathbf{y}$ and 00 otherwise. Finally, define $[w]_{3}=1$ if $w^{R}$ is also a factor of $\mathbf{y}$ but $w$ is not a palindrome, and 00 otherwise. Notice that these three cases are disjoint and subsume all possibilities for factors of $\mathbf{y}$ (also recall the decomposition at the beginning of the section). We can extend this notation to sets by defining $[S]_{i}=\sum_{w\in S}[w]_{i}$ for $i\in\{1,2,3\}$. Define $[w]=[w]_{1}+[w]_{2}+[w]_{3}/2$ and similarly for $[S]$. From Equalities (5), we know that
@@
Now suppose $w$ is not exceptional. There are three cases to consider.
@@
Numerical experiments concerning the growth of the reflection complexity have led us to formulate Conjectures 2427 below. We leave these conjectures as open problems.
Conjecture 24.
Conjecture 25.
Let $\mathbf{x}$ be a sequence. Then $r_{\mathbf{x}}(n)-1\leq r_{\mathbf{x}}(n+1)$ for all $n\geq 0$.
Note that we can have equality for infinitely many values of $n$—see Example 16 above.
Conjecture 26.
It can be shown that Conjecture 26 holds for the Thue–Morse, period-doubling, Golay–Shapiro, second-bit, paperfolding, Stewart choral, Baum-Sweet, Chacon, and Mephisto-Waltz sequences. (Also see Corollary 30.)
Conjecture 27.
Actually we prove below weaker forms of Conjectures 25 and 26 for reversal-closed sequences, and weaker forms of Conjectures 24 and 25 for sequences without long palindromes. Also we can prove that Conjecture 27 holds for primitive morphic sequence.
Theorem 28.
Let $\mathbf{x}$ be a reversal-closed sequence. Then, for all $n\geq 0$, we have $r_{\mathbf{x}}(n)-1\leq r_{\mathbf{x}}(n+1)$. If, in addition, $\mathbf{x}$ has at most linear factor complexity, then $r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n)$ is bounded.
Proof.
Let $\mathbf{x}$ be a reversal-closed sequence. By Remark 10, $\mathbf{x}$ is recurrent. Thus, putting together Theorems 17(b) and 9(b) gives, for all $n\geq 0$,
$2(r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n))=\rho_{\mathbf{x}}(n+1)+\operatorname{Pal}_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n)-\operatorname{Pal}_{\mathbf{x}}(n).$ (6)
Equality (6) implies $22(r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n))\geq\operatorname{Pal}_{\mathbf{x}}(n+1)-$ 2, hence we obtain $r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n)\geq-1$ as desired.
If, in addition, the factor complexity of $\mathbf{x}$ is at most linear, then Equality (6) gives
$2|r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n)|\leq|\rho_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n)|+\operatorname{Pal}_{\mathbf{x}}(n+1)+\operatorname{Pal}_{\mathbf{x}}(n).$
But $|\rho_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n)|$ is bounded (see [27]) and $\operatorname{Pal}_{\mathbf{x}}$ is also bounded (see [4, Theorem 12] or use Theorem 17(b) above). ∎
@@
$\rho_{\mathbf{x}}(n+2)+\operatorname{Unr}_{\mathbf{x}}(n+2)=\rho_{\mathbf{x}}(n+1)+\operatorname{Unr}_{\mathbf{x}}(n+1)=\rho_{\mathbf{x}}(n)+\operatorname{Unr}_{\mathbf{x}}(n).$
Hence $\rho_{\mathbf{x}}(n+1)=\rho_{\mathbf{x}}(n)$, which implies that $\mathbf{x}$ is eventually periodic from Theorem 5, as desired. ∎
Actually, Theorems 11, 28 and 29 imply the following corollary.
Corollary 30.
Conjectures 25 and 26 hold if $\mathbf{x}$ is uniformly recurrent.
Proof.
Let $\mathbf{x}$ be a uniformly recurrent sequence. Then, from Theorem 11, we have that $\mathbf{x}$ is either reversal-closed and $\rho_{\mathbf{x}}=\operatorname{Ref}_{\mathbf{x}}$, so that $r_{\mathbf{x}}=\frac{1}{2}(\rho_{\mathbf{x}}+\operatorname{Pal}_{\mathbf{x}})$, or else that there exists $n_{0}$ such that $\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)$ for all $n\geq n_{0}$, which implies $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$ for all $n\geq n_{0}$.
In the first case, the claim is proved by using Theorem 28, and that, if, in addition, $\mathbf{x}$ is automatic, then both sequences $(\rho_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n))_{n\geq 0}$ and $(\operatorname{Pal}_{\mathbf{x}}(n))_{n\geq 0}$ are automatic (see [26, Theorem 4.3] and [26, Theorem 4.8] respectively). Hence the sequence $(\operatorname{Pal}_{\mathbf{x}}(n+1)-\operatorname{Pal}_{\mathbf{x}}(n))_{n\geq 0}$ is also automatic. The proofs extend easily to generalized automatic sequences.
In the second case, inspired by the proof of Theorem 29, we note that there is an integer $n_{0}$ such that $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)$ for all $n\geq n_{0}$. But $\operatorname{Unr}_{\mathbf{x}}$ is clearly non-decreasing, which gives the property of Conjecture 25. If, in addition, $\mathbf{x}$ has at most linear complexity, we know that $\rho_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n)$ is bounded (see [27]), hence $r_{\mathbf{x}}(n+1)-r_{\mathbf{x}}(n)$ is bounded for $n$ large enough, hence for all $n$. The (generalized) automatic property is proved by using, as above, that $(\rho_{\mathbf{x}}(n+1)-\rho_{\mathbf{x}}(n))_{n\geq 0}$ is (generalized) automatic. ∎
In the same vein, Theorem11 and Corollary 21 imply the following corollary.

### v4

@@ 3 General results
Given a sequence $\mathbf{x}$, we can decompose its factor complexity function $\rho_{\mathbf{x}}$ and its reflection complexity function $r_{\mathbf{x}}$ by using the following functions: for all $n\geq 0$, we let
@@
    $\operatorname{Unr}_{\mathbf{x}}(n)$ denote the number of “unreflected” length-$n$ factors $w$ of $\mathbf{x}$ such that $w^{R}$ is not a factor of $\mathbf{x}$;
@@
    $\operatorname{Ref}_{\mathbf{x}}(n)$ denote the number of “reflected” length-$n$ factors $w$ of $\mathbf{x}$ such that $w^{R}$ is also a factor of $\mathbf{x}$; and
@@
    $\operatorname{Pal}_{\mathbf{x}}(n)$ denote the number of length-$n$ palindrome factors $w$ of $\mathbf{x}$ (i.e., the palindrome complexity function of $\mathbf{x}$ [4]).
@@
$\begin{array}[]{l}\rho_{\mathbf{x}}=\operatorname{Unr}_{\mathbf{x}}+\operatorname{Ref}_{\mathbf{x}},\\ r_{\mathbf{x}}=\operatorname{Unr}_{\mathbf{x}}+\dfrac{1}{2}(\operatorname{Ref}_{\mathbf{x}}-\operatorname{Pal}_{\mathbf{x}})+\operatorname{Pal}_{\mathbf{x}}.\end{array}$ (5)
@@
Let $\mathbf{f}=01001010\cdots$ denote the Fibonacci sequence, which is the fixed point of $0\mapsto 01,1\mapsto 0$. Its length-$5$ factors are $u_{1}=01001$, $u_{2}=10010$, $u_{3}=00101$, $u_{4}=01010$, $u_{5}=10100$, and $u_{6}=00100$. Observe that $u_{1},u_{2},u_{3},u_{5}$ are reflected (second type), and $u_{4}$ and $u_{6}$ are palindromes (second and third types). Furthermore $\mathbf{f}$ has no unreflected factor, because Sturmian words are reversal-closed; see [47, Theorem 4, p. 77]. We therefore obtain $r_{\mathbf{f}}(6)=0+\frac{1}{2}(6-2)+2=4$.
The relationships among the complexity functions $\operatorname{Unr}_{\mathbf{x}}$, $\operatorname{Ref}_{\mathbf{x}}$, $\rho_{\mathbf{x}}$, $r_{\mathbf{x}}$, and $\operatorname{Pal}_{\mathbf{x}}$ motivates the study of the combinations of these functions indicated in Equalities (5). This is illustrated below.
@@
This lemma implies the following bounds on the ratio $r/\rho$.
@@
For a sequence $\mathbf{x}$ and for all $n\geq 0$, we have
@@
Furthermore, the equality cases are as follows:
@@
It is known that if a sequence $\mathbf{x}$ is reversal-closed, then $\mathbf{x}$ is recurrent: it suffices to adapt the proof of [43, Proposition 1, p. 176], as indicated in [22]. Also note that if a sequence $\mathbf{x}$ is uniformly recurrent and contains infinitely many distinct palindromes, then $\mathbf{x}$ is reversal-closed [15, Theorem 3.2].
@@
If $\mathbf{x}$ is reversal-closed, then $r_{\mathbf{x}}=\frac{1}{2}(\rho_{\mathbf{x}}+\operatorname{Pal}_{\mathbf{x}})$ from Theorem 9(b) above. Now suppose that $\mathbf{x}$ has an unreflected factor $w$. Since $\mathbf{x}$ is uniformly recurrent, every sufficiently long factor of $\mathbf{x}$ contains $w$ as a factor, which implies that this long factor itself is unreflected. This exactly says that $\operatorname{Ref}_{\mathbf{x}}(n)=0$ for $n$ large enough (and in particular $\operatorname{Pal}_{\mathbf{x}}(n)=0$ for $n$ large enough). This implies from Equalities (5) that, for $n$ large enough, $\rho_{\mathbf{x}}(n)=\operatorname{Unr}_{\mathbf{x}}(n)$, and so $r_{\mathbf{x}}(n)=\rho_{\mathbf{x}}(n)$. ∎
@@
Consider the sequence $\mathbf{x}$ on $\{0,1\}$ whose $n$th prefix $x_{n}$ is given recursively as follows: $x_{0}=01$ and $x_{n+1}=x_{n}01x_{n}^{R}$ for all $n\geq 0$. See [19, Section 3] or [15, Example 3.1]. The sequence $\mathbf{x}$ is uniformly recurrent, reversal-closed, $2$-automatic, and accepted by a DFAO of 6 states (see, e.g., [6]), and contains only a finite number of palindromes. Furthermore, for all sufficiently large $n$, we have $r_{\mathbf{x}}(n)=\dfrac{1}{2}\rho_{\mathbf{x}}(n)$.
@@
It is also possible to construct an aperiodic automatic sequence where the only palindromes are of length 1, but there are reflected factors of each length $>1$. Let $g_{n}$ be the prefix of length $22^{n}-$ 2 of $(012)^{\omega}$. Then an example of an automatic sequence satisfying the desired properties is
@@
where [80, Theorem 1] is required (observe that we intertwine the sequences $(3456)^{\omega}$ and $g_{1}g_{2}^{R}g_{3}g_{4}^{R}\cdots$ to build $\mathbf{x}$).
@@
We also provide a construction of an automatic sequence $\mathbf{x}$ such that $r_{\mathbf{x}}(n+1)<r_{\mathbf{x}}(n)$ for all odd $n\geq 3$. In particular, let $\mathbf{x}$ denote the sequence given by applying the coding $a,b,d\mapsto 1$ and $c\mapsto 0$ to the fixed point, starting with $a$, of the morphism defined by $a\mapsto ab$, $b\mapsto cd$, $c\mapsto cd$, and $d\mapsto bb$. This gives us sequence [92, A039982] in the OEIS. Computing the reflection complexity of $\mathbf{x}$ (e.g., using Walnut) gives that
@@
Actually we even have that $r_{\mathbf{x}}(n+1)=r_{\mathbf{x}}(n)-1$ for all odd $n\geq 3$. See Theorem 17.
With extra hypotheses on a sequence $\mathbf{x}$, we can give more precise results in comparing the respective growths of reflection and factor complexities. We will need Theorem 17 below. Note that Part (b) of this theorem was originally stated for uniformly recurrent sequences: see [14, Theorem 1.2]. However, its proof only requires the sequences to be recurrent (see [15, p. 449] and also [24, Footnote, p. 493]). Furthermore we have seen that a reversal-closed sequence must be recurrent (see Remark 10). Thus we can state the theorem as follows (also see Theorem 11).
@@
We deduce the following results from Theorem 17.
@@
Using the first inequality in Theorem 9, the statement in Theorem 9(b), and Theorem 17(b), we have
@@
for all $n\geq 0$. The desired inequalities follow. ∎
Proposition 20.
Proof.
It is enough to combine Theorems 17(b) and 9(b). ∎
@@
for all $n\geq 1$. The statement follows from this inequality and Theorem 9. ∎
@@
when $n$ tends to infinity. In particular, this equivalence holds if $\mathbf{x}$ is non-eventually periodic, reversal-closed, and morphic.
@@
Let $\mathbf{x}$ be a non-eventually periodic and reversal-closed sequence. If $\rho_{\mathbf{x}}(n+1)\sim\rho_{\mathbf{x}}(n)$, then, from Theorem 19, we obtain that $r_{\mathbf{x}}(n)\sim\frac{1}{2}\rho_{\mathbf{x}}(n)$ when $n$ tends to infinity. Now, if $\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}=o(n)$, we obtain, from Theorem 21, and using the fact that $\rho_{\mathbf{x}}$ is non-decreasing, that
@@
Now suppose that, in addition, the sequence $\mathbf{x}$ is morphic. We know that either $\rho_{\mathbf{x}}(n)=\Theta(n^{2})$ or $\rho_{\mathbf{x}}(n)=O(n^{3/2})$ (see [45, 44]). In the first case, then $\frac{\rho_{\mathbf{x}}(2n)}{\rho_{\mathbf{x}}(n)}$ is bounded, and hence $o(n)$. If $\rho_{\mathbf{x}}(n)=O(n^{3/2})$, since ${\mathbf{x}}$ is not eventually periodic (hence $\rho_{\mathbf{x}}(n)\geq n+1$), there exists a constant $C>0$ such that
@@
The upper bound in Theorem 21 raises questions about the growth properties of the function $r_{\mathbf{x}}$ more generally, apart from the case where the set of factors of $\mathbf{x}$ satisfies the hypotheses of Theorem 21. This leads us to the growth property in Theorem 23 below.
@@
For a factor $w$ of $\mathbf{y}$, define $[w]_{1}=1$ if $w$ is a palindrome, and 00 otherwise. Similarly, $[w]_{2}=1$ if $w^{R}$ is not a factor of $\mathbf{y}$ and 00 otherwise. Finally, define $[w]_{3}=1$ if $w^{R}$ is also a factor of $\mathbf{y}$ but $w$ is not a palindrome, and 00 otherwise. Notice that these three cases are disjoint and subsume all possibilities for factors of $\mathbf{y}$ (also recall the decomposition at the beginning of the section). We extend this notation to sets by defining $[S]_{i}=\sum_{w\in S}[w]_{i}$ for $i\in\{1,2,3\}$. Define $[w]=[w]_{1}+[w]_{2}+[w]_{3}/2$ and similarly for $[S]$. From Equalities (5), we know that
@@
Now suppose $w$ is not exceptional. There are three cases to consider, noting that when $n=1$, we have that $[w]_{1}=1$ and $[w]_{2}=0$.
@@
With regard to the conjectures listed at the end of this section, we can prove a weaker form of Conjecture 28 for reversal-closed sequences, and a weaker form of Conjecture 27 for sequences without long palindromes. Also we can prove that Conjecture 29 holds for primitive morphic sequence.
@@
$\rho_{\mathbf{x}}(n+2)+\operatorname{Unr}_{\mathbf{x}}(n+2)=\rho_{\mathbf{x}}(n+1)+\operatorname{Unr}_{\mathbf{x}}(n+1)=\rho_{\mathbf{x}}(n)+\operatorname{Unr}_{\mathbf{x}}(n).$ (7)
Hence $\rho_{\mathbf{x}}(n+1)=\rho_{\mathbf{x}}(n)$, which implies that $\mathbf{x}$ is eventually periodic from Theorem 5. ∎
Numerical experiments concerning the growth of the reflection complexity have led us to formulate Conjectures 2729 below. Actually, Theorems 11 and 25 can be used to show that Conjecture 28 holds if $\mathbf{x}$ is uniformly recurrent (also see Theorem 30 below). In the same vein, Theorem 11 and Corollary 22 imply the following corollary.